OpenStudy (anonymous):

Which of the following equations does not require two steps to solve? -32+k=-26 -3b-22=7 5w-19=-15 10-p=2

OpenStudy (anonymous):

second and third requires 3 steps i think

OpenStudy (anonymous):

But, Which one for shure?

OpenStudy (phi):

when you solve all 4 equations, the first step is to move the constant term from the left side to the right side. then you would divide both sides of the equation by the coefficient of the variable to complete the process. however, note that the coefficient of k in the first equation is 1, so you do not need to do this step

OpenStudy (anonymous):

I dont understand....

OpenStudy (anonymous):

can you answer both? or just one? cuz i think it's both second n third.. in the second one, we have to exchange -3b and 7, so that the b would be positive, and then evaluate -7 -22 = -29 then -29 divided by 3 so it's 3 steps the same as third one

OpenStudy (anonymous):

I can only answer one.. lol this is confusing mee!

OpenStudy (phi):

to solve the first equation -32+k=-26 you would add 32 to both sides: -32+k= -26 +32 = +32 ------------- k= + 6 That is a "one step" problem

OpenStudy (anonymous):

Do you know which one is the answer?

OpenStudy (phi):

the question is asking Which of the following equations does not require two steps to solve?

OpenStudy (anonymous):

I know tht

OpenStudy (anonymous):

yea i think you can answer both

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

thanks i got tht tho! :)